# Is That Your Final Answer? Test-Time Scaling Improves Selective Question Answering

### v1

@@ Abstract
Scaling the test-time compute of large language models has demonstrated impressive performance on reasoning benchmarks. However, existing evaluations of test-time scaling make the strong assumption that a reasoning system should always give an answer to any question provided. This overlooks concerns about whether a model is confident in its answer, and whether it is appropriate to always provide a response. To address these concerns, we extract confidence scores during reasoning for thresholding model responses. We find that increasing compute budget at inference time not only helps models answer more questions correctly, but also increases confidence in correct responses. We then extend the current paradigm of zero-risk responses during evaluation by considering settings with non-zero levels of response risk, and suggest a recipe for reporting evaluations under these settings.
@@
Figure 1: DeepSeek R1-32B’s accuracy is a function of compute budget and confidence threshold. Increased confidence thresholds generally yield increased accuracy at the cost of response rate, while increased compute budgets sometimes decrease accuracy while increasing response rate. The vertical axis measures the accuracy of answered questions at a compute budget and confidence threshold. Color indicates the proportion of questions that are answered; in redder regions, the model is more likely to answer, whereas in bluer regions the model is less likely to answer. We treat the case where the model never answers as accuracy 0.
Figure 2: Confidence thresholds on test-time scaling. (left) When the logit threshold is 0, the model answers 100% of questions. This is the only performance curve that is reported by test-time scaling research. (center) At a moderate threshold, more frequent absentions allow higher response accuracy. (right) At a high threshold, small amounts of test-time compute deliver very high accuracy, while test-time scaling provides more answers at the cost of answer accuracy. We treat the decision to never answer as yielding accuracy 0.
@@ 1 Introduction
Many question answering settings associate incorrect answers with measurable costs, ranging from low-risk responses found in game shows (Ferrucci et al., 2010) to high-stakes responses that can alter people’s lives (Northpointe, 2017). Selective question answering addresses these challenges by allowing a model to refrain from answering questions which it might answer incorrectly Kamath et al. (2020). This requires a selection function, which considers risk tolerance, coverage goals, and candidate answer confidence to decide whether a prediction should be given Geifman and El-Yaniv (2017). Knowing when not to answer is a critical for systems to collaborate effectively with humans Verma et al. (2023), especially for test-time scaling systems that must constantly decide between refusing to answer and expending further compute to search for a possible solution.
To help address this issue, we evaluate test-time scaling models using a simple class of selection functions that reject questions if a model is not confident in its answer after expending its compute budget. We evaluate these systems at different compute budgets, showing a new aspect of model performance that answer accuracy alone struggles to measure. We suggest a class of utility functions that represent various levels of error risk to empirically measure the performance of these systems in settings where incorrect answers are penalized. Evaluation in these settings shows how compute scaling affects confidence in existing systems. Based on these insights, we propose a standard method for measuring model performance in settings with non-zero response risk. In summary we:
@@
  • Introduce evaluation settings that penalize incorrect answers and allow abstentions to help holistically models capable of scaling test-time compute. (Section 4)
@@ 2 Methods
We explore how increasing compute budgets affects a model’s performance on QA tasks at different confidence thresholds. The choice of a budget and threshold is a test-time decision. We describe methods to quantify the two factors below:
Compute Budget refers to the amount of compute expended by the model at inference time. In all cases, we quantify a model’s budget by counting the number of tokens in its reasoning trace. We use methods proposed by Muennighoff et al. (2025) to strictly enforce compute budgets. Specifically, we ignore any predicted end-of-thinking delimiters and instead append the token “Wait” if a model attempts to end its reasoning trace before reaching the budget, and we force decode the end-of-thinking delimiter once the budget is reached.
Confidence Threshold refers to the uncertainty of the model in its decoded answer. We quantify a model’s confidence as the sum of the log-probabilities corresponding to the answer tokens. [^1] Every answer in our dataset is a 3-digit number between 000 and 999, so consists of the same number of tokens. For a confidence threshold, our selection function Geifman and El-Yaniv (2017) only accepts answers that the model delivers with confidence greater than its threshold, abstaining otherwise.
@@ 3 Experiments
Experimental Setup
We evaluate Deepseek-R1-32B (DeepSeek-AI et al., 2025) and s1 (Muennighoff et al., 2025) due to their exhibited test-time scaling capabilities and open-weight checkpoints, and choose AIME24 as our evaluation dataset. The dataset contains 30 hard math problems on which performance substantially benefits from larger compute budgets, making it a popular benchmark for evaluating test-time scaling. We test the set of confidence thresholds $\{0.0,0.5,0.95\}$ across compute budgets within the range $[500,8000]$, incrementing by 100 tokens. For a given budget and threshold, we report the accuracy of answered questions. As the number of answered questions differs across confidence thresholds, we note that accuracies are not directly comparable between models and compute budgets.
Figure 3: Test-time scaling improves confidence in correct answers. Each dot represents R1 32B’s confidence in an answer after spending a fixed amount of compute. Indigo series are correct answers, while orange series are incorrect. Note that individual dots may turn from orange to indigo if the model changes its prediction after thinking longer. See Figure 7 in Appendix B for s1-32B.
Results
Figure 2 compares the accuracy of answers provided by R1-32B and s1-32B at different test-time compute budgets. When the confidence threshold is 0, models answer every question, so accuracy increases consistently with compute budget. We observe that these subplots are slices of a surface parameterized by compute budget and confidence threshold, shown in Figure 1. While higher confidence thresholds prevent the model from answering at low budgets, scaling compute at high thresholds delivers a larger volume of accurate answers. At higher confidence thresholds, increased compute budget can actually decrease answer accuracy. This decrease in accuracy of yielded answers does not necessarily reflect decreased performance at higher budgets, but instead the increase in the total number of questions answered.
To investigate whether excessive thinking harms accuracy drops by pushing models to abandon correct answers, we plot how a model’s confidence in individual answers moves over time. Figure 3 shows the answer confidences given by R1-32B at varying compute budgets, colored according to their correctness, with a curve fit to the distribution. We note that as compute budget increases, the average confidence of its correct answers increases even as additional correct answers are discovered.
@@ 4 Utility
Motivation
When refusal to answer is an option, accuracy can be trivially optimized by a system that answers extremely infrequently. Thus, a useful metric must capture both the accuracy of answers provided and the system’s propensity to provide answers. Many real world scenarios reward correct answers, but incur measurable costs for incorrect answers. We show our results involving confidence thresholds can be adapted to these settings: Given a model $\mathcal{M}$ and an instance $x$ of a task $t$, we define a utility function $f$ to be $f(\mathcal{M},x)=\mathbbm{1}_{c}+0\mathbbm{1}_{a}+r_{t}\mathbbm{1}_{i}$ where $(\mathbbm{1}_{c},\mathbbm{1}_{a},\mathbbm{1}_{i})$ are mutually exclusive indicators for the predicted answer being (correct, abstained, incorrect) and $r_{t}$ is some task-specific cost of incorrect answers. We can assume the reward for correct answers is 1 without loss of generality due to scaling. While there exist scenarios where refusing to answer also incurs a cost, this paper will only discuss the consequences when no extra cost is incurred; the conclusions we draw can be extended to these cases.
Problem Scenarios
@@
  • Jeopardy Odds [^2] Inspired by the wagers made in the ‘Final Jeopardy’ stage ($r_{t}=-1$): The cost of an incorrect answer is equal to the reward for a correct answer. In these scenarios, no answer at all is preferable to an incorrect answer.
@@
Figure 5: Jeopardy utility scales differently across models and thresholds. Performance of s1-32B and R1-32B in the Jeopardy odds setting under different confidence thresholds. While s1 is competitive in the case when threshold is 0, a higher threshold shows R1’s superior scaling performance.
Results
We keep the same experimental setup as described in Section 3. Rather than reporting accuracy, we instead report the utility in the three scenarios above, shown in Figure 4. We focus on the Jeopardy setting because it highlights why a system might choose not to answer; results in the other settings are in Appendix B.
@@ 5 Related Work
As scaling training compute has become prohibitively expensive Hoffmann et al. (2022), models that scale performance with test-time compute have become a new frontier Snell et al. (2024); Wu et al. (2024). These methods have delivered state-of-the-art results on hard reasoning tasks using lengthy chains of thought DeepSeek-AI et al. (2025); Muennighoff et al. (2025). Current work in this space optimizes for question answering tasks which do not penalize incorrectness, ignoring settings that favor refusal over wrong answers Ferrucci et al. (2010); Rajpurkar et al. (2018); Kamath et al. (2020). We draw motivation from methods for cost-sensitive learning Mienye and Sun (2021) and selective classification Geifman and El-Yaniv (2017), which navigate penalties for failure. These setting reward confidence calibration, which can be critical for effective collaboration with human experts Verma et al. (2023). We are the first to investigate how serialized test-time compute helps models identify when they should not answer. We include additional materials in Appendix A for further reading on these areas.
@@ 6 Conclusion
The selection function we implement is based entirely on the likelihood that a large language model assigns a series of tokens after thinking, which is not necessarily the optimal method for model confidence estimation. We choose the AIME dataset to demonstrate the value of our proof because it displays the most consistent performance improvements from test-time scaling Muennighoff et al. (2025); We evaluate a fixed set of three odds levels, but recognize that this does not cover the full space of real-world utility functions. Furthermore, we do not consider how compute costs might be incorporated in the model’s utility function, which could encourage increased energy consumption. Finally, we recognize that by evaluating only on English mathematics questions and answers, we may miss model capabilities or weaknesses in lower-resource languages.
@@ Appendix A Appendix
A.1 Supplementary Background
Test-time Scaling
Many methods for scaling test-time comptue have been explored. These include searching over possible generations (Wang et al., 2024; Snell et al., 2024), tool use (Qin et al., 2024; Qu et al., 2025), making gradient updates at inference time Akyürek et al. (2024); Li et al. (2025), and simply fine-tuning on longer chains of thought (DeepSeek-AI et al., 2025; Muennighoff et al., 2025). Our work considers models fine-tuned on extremely long reasoning chains, and augments them with the ability to refuse to answer questions where they lack confidence. Concurrent work fine-tunes similar models to consider the difficulty of a question at when allocating compute test-time, but do not report the effect their method has on their models’ confidence Arora and Zanette (2025). In contrast, we show how model confidence scales with test-time compute, and demonstrate its value for optimizing performance in settings that allow refusal to answer.
Answer Refusal
Refusing to answer is an important choice in many prior works on question answering. SQuAD 2.0 included this feature by asking questions which have no answer, although they treat abstaining from answering as a correct answer in these unanswerable cases Rajpurkar et al. (2018). Game-show based research efforts use an approach more closely aligned with ours, which penalizes systems for answering incorrectly to encourage abstentions when a system cannot develop sufficiently high confidence Ferrucci et al. (2010); Ferrucci (2012); Boyd-Graber and Börschinger (2020); Rodriguez et al. (2021). Related to these quiz game based setting is research into selective classification, which evaluates models performance across the coverage-accuracy curve, rather than at single point Geifman and El-Yaniv (2017). These approaches can be useful for avoiding costly errors in high-pressure domains Khan et al. (2018) or under distribution shift Ren et al. (2023), or when designing systems that defer to expert humans when it lacks confidence that its input will be helpful Mozannar and Sontag (2020). Related research in language modeling has focused on calibrating prediction confidence to align with human judgments Jiang et al. (2024). Most recently, research has investigated training language models to refuse to answer Cao (2024). However, this does not ask how this behavior should be measured effectively in sequential test-time compute models, which might find a confident answer given higher compute budgets.
A.2 Implementation Details
We use widely available open-source libraries to run our experiments, including HuggingFace Transformers Wolf et al. (2020) and vLLM Kwon et al. (2023) for language model inference, and the Language Model Evaluation Harness Gao et al. (2024) to sample reasoning chains for the AIME problems at temperature 0. In particular, we use the variant of this library released by Muennighoff et al. (2025), and run a subset of the experiments that they run. We run experiments on a node with 4 H100 GPUs for approximately 4 hours.
Appendix B Supplemental Figures
Figure 6: s1-32B’s answer accuracy is a function of compute budget and confidence threshold. Increased confidence thresholds generally yield increased accuracy at the cost of response rate, while increased compute budgets sometimes decrease accuracy while increasing response rate. The vertical axis indicates the accuracy for answered questions at a compute budget and logit threshold. The color indicates the proportion of questions that are answered; in redder regions, the model is more likely to answer, whereas in bluer regions the model is less likely to answer. We treat the decision to never answer as accuracy 0.
Figure 7: Test-time scaling improves confidence in correct answers. Each dot represents s1-32B’s confidence in an answer after spending a fixed amount of compute. Indigo series indicate correct answers, while orange series are incorrect. Note that individual dots may switch colors if the model changes its prediction after thinking longer. s1-32B does not separate its correct answers from its incorrect answers as effectively as R1-32B. See Figure 3 for R1-32B.
Figure 8: Utility Surface of s1-32B for Jeopardy. The vertical axis indicates performance in the Jeopardy setting at different compute budgets and confidence thresholds. The color indicates the proportion of questions that are answered; in redder regions, the model is more likely answer, whereas in bluer regions the model is less likely to answer. The horizontal plane divides positive and negative utility regions of the operating curve. The checkered lines indicate the threshold slices that we compare against R1-32B in Figure 5. Note the relatively lower volume above the break-even point of 0 corresponds to s1-32B’s broadly inferior performance at these odds.

### v2

@@ Abstract
Scaling the test-time compute of large language models has demonstrated impressive performance on reasoning benchmarks. However, existing evaluations of test-time scaling make the strong assumption that a reasoning system should always give an answer to any question provided. This overlooks concerns about whether a model is confident in its answer, and whether it is appropriate to always provide a response. To address these concerns, we extract confidence scores during reasoning for thresholding model responses. We find that increasing compute budget at inference time not only helps models answer more questions correctly, but also increases confidence in correct responses. We then extend the current paradigm of zero-risk responses during evaluation by considering settings with non-zero levels of response risk, and suggest a recipe for reporting evaluations under these settings. [^1] Code released at https://github.com/wjurayj/final_answer
@@
Figure 1: DeepSeek R1-32B’s accuracy is a function of compute budget and confidence threshold. Increased confidence thresholds generally yield increased accuracy at the cost of response rate, while increased compute budgets sometimes decrease accuracy as they increase response rate. The vertical axis measures the accuracy of answered questions at a compute budget and confidence threshold. Color indicates the proportion of questions that are answered; in redder regions, the model is more likely to answer, whereas in bluer regions the model is less likely to answer.
Figure 2: Confidence thresholds on test-time scaling. (left) When the confidence threshold is 0, the model answers 100% of questions. This is the only performance curve that is reported by test-time scaling research. (center) At a moderate threshold, more frequent absentions allow higher response accuracy. (right) At a high threshold, small amounts of test-time compute deliver very high accuracy at low answer rates, while test-time scaling provides more answers at the cost of answer accuracy.
@@ 1 Introduction
Many question answering settings associate incorrect answers with measurable costs, ranging from low-risk responses found in game shows (Ferrucci et al., 2010) to high-stakes responses that can alter people’s lives (Northpointe, 2017). Selective question answering addresses these challenges by allowing a model to refrain from answering questions which it might answer incorrectly Kamath et al. (2020). This requires a selection function, which considers risk tolerance, coverage goals, and candidate answer confidence to decide whether a prediction should be given Geifman and El-Yaniv (2017). Knowing when not to answer is a critical quality for systems to collaborate effectively with humans Verma et al. (2023), especially for test-time scaling systems that must constantly decide between refusing to answer and expending further compute to search for a possible solution.
To help address this issue, we evaluate test-time scaling models using a simple class of selection functions that reject questions if a model is not confident in its answer after expending its compute budget. We evaluate these systems at different compute budgets, showing a new axis of model performance that answer accuracy alone struggles to measure. We suggest a class of utility functions that represent various levels of error risk to empirically measure the performance of these systems in settings where incorrect answers are penalized. Evaluation in these settings shows how compute scaling affects confidence in existing systems. Based on these insights, we propose a standard method for measuring model performance in settings with non-zero response risk. In summary we:
@@
  • Introduce evaluation settings that penalize incorrect answers and allow abstentions to help holistically evaluate models capable of scaling test-time compute. (Section 4)
@@ 2 Methods
We explore how increasing compute budgets affects a model’s performance on question answering tasks at different confidence thresholds. The choice of a budget and threshold is a test-time decision. We describe methods to quantify the two factors below:
Compute Budget
refers to the amount of compute expended by the model at inference time. In all cases, we quantify a model’s budget by counting the number of tokens in its reasoning trace. We use methods proposed by Muennighoff et al. (2025) to strictly enforce compute budgets. Specifically, we ignore any predicted end-of-thinking delimiters and instead append the token “Wait” if a model attempts to end its reasoning trace before reaching the budget, and we force decode the end-of-thinking delimiter once the budget is reached.
Confidence Threshold
refers to the uncertainty of the model in its decoded answer. We quantify a model’s confidence as the sum of the log-probabilities corresponding to the answer tokens. [^2] Every answer in our dataset is a 3-digit number between 000 and 999, so consists of the same number of tokens. For a confidence threshold, our selection function Geifman and El-Yaniv (2017) only accepts answers that the model delivers with confidence greater than its threshold, abstaining otherwise.
Figure 3: Test-time scaling improves confidence in correct answers on R1-32B (left) and S1-32B (right). Each dot represents the model’s confidence in an answer after spending a fixed amount of compute. Indigo series are correct answers, while orange series are incorrect. Dotted lines plot confidence trajectories for 10 randomly selected questions, emphasizing how confidence in correct versus incorrect answers changes with test-time compute. Note that individual answers may turn from orange to indigo if the model changes its prediction after thinking longer.
@@ 3 Experiments
3.1 Experimental Setup
We evaluate Deepseek-R1-32B (DeepSeek-AI et al., 2025) and s1 (Muennighoff et al., 2025) due to their exhibited test-time scaling capabilities and open-weight checkpoints, and choose AIME 2024 and 2025 as our primary evaluation dataset. This dataset contains 60 hard math problems on which performance substantially benefits from larger compute budgets, making it a popular benchmark for evaluating test-time scaling. Additional experiments on GPQA Rein et al. (2024) are included in Appendix B. We test the set of confidence thresholds $\{0.0,0.5,0.95\}$ across compute budgets within the range $[500,8000]$, incrementing by 100 tokens. For a given budget and threshold, we report the accuracy of answered questions, treating never answering as yielding accuracy 0. As the number of answered questions differs across confidence thresholds, we note that accuracies are not directly comparable between models and compute budgets.
We use widely available open-source libraries to run our experiments, including HuggingFace Transformers Wolf et al. (2020) and vLLM Kwon et al. (2023) for language model inference, and the Language Model Evaluation Harness Gao et al. (2024) to sample reasoning chains at temperature 0 and 32-bit precision. In particular, we use the variant of this library released by Muennighoff et al. (2025), and run a subset of the experiments that they run. We run experiments on 4 H100 GPUs.
3.2 Results
Figure 2 compares the accuracy of answers provided by R1-32B and S1-32B at different test-time compute budgets. When the confidence threshold is 0, models answer every question, so accuracy increases consistently with compute budget. We observe that these subplots are slices of a surface parameterized by compute budget and confidence threshold, shown in Figure 1. While higher confidence thresholds prevent the model from answering at low budgets, scaling compute at high thresholds delivers a larger volume of accurate answers. However, at higher confidence thresholds increased compute budget can actually decrease answer accuracy. This decrease in accuracy of yielded answers does not necessarily reflect decreased performance at higher budgets, but instead that the additional questions answered are less likely to be correct than those answered at lower budgets.
To investigate whether excessive thinking harms accuracy drops by pushing models to abandon correct answers, we plot how a model’s confidence in individual answers moves over time. Figure 3 shows the answer confidences given by both models at varying compute budgets, colored according to their correctness, with a cubic curve fit to the distribution. We note that as compute budget increases, the average confidence of its correct answers increases even as additional correct answers are discovered. Notably, this is not a universal property of test-time scaling models: S1-32B does not separate its correct answers from its incorrect answers as well as R1-32B.
@@ 4 Utility
4.1 Motivation
When refusal to answer is an option, accuracy can be trivially optimized by a system that answers extremely infrequently. Thus, a useful metric must capture both the accuracy of answers provided and the system’s propensity to provide answers. Many real world scenarios reward correct answers, but incur measurable costs for incorrect answers. We show our results involving confidence thresholds can be adapted to these settings.
Given a model $\mathcal{M}$ and an instance $x$ of a task $t$, we define a utility function $f$ to be
$f(\mathcal{M},x)=\begin{cases}1&\textrm{ $\mathcal{M}$ answers $x$ correctly}\\ 0&\textrm{ $\mathcal{M}$ abstains from answering $x$ }\\ r_{t}&\textrm{ $\mathcal{M}$ answers $x$ incorrectly}\end{cases}$
We can assume the reward for correct answers is 1 without loss of generality due to scaling. While there exist scenarios where refusing to answer also incurs a cost, this paper will only discuss the consequences when no extra cost is incurred; the conclusions we draw can be extended to these cases.
4.2 Problem Scenarios
@@
  • Jeopardy Odds [^3] Inspired by the wagers made in the game show’s ‘Final Jeopardy’ stage ($r_{t}=-1$): The cost of an incorrect answer is equal to the reward for a correct answer. In these scenarios, no answer at all is preferable to an incorrect answer.
@@
4.3 Results
We keep the same experimental setup as described in Section 3.1. Rather than reporting accuracy, we instead report the utility in the three scenarios above, shown in Figure 4. We focus on the Jeopardy setting because it highlights why a system might choose not to answer; results in the other settings are in Appendix A.
@@ 5 Related Work
As scaling training compute has become prohibitively expensive Hoffmann et al. (2022), models that scale performance with test-time compute have become a new frontier Snell et al. (2024); Wu et al. (2025). These methods have delivered state-of-the-art results on hard reasoning tasks using lengthy chains of thought DeepSeek-AI et al. (2025); Muennighoff et al. (2025). Current work in this space optimizes for question answering tasks which do not penalize incorrectness, ignoring settings that favor refusal over wrong answers Ferrucci et al. (2010); Rajpurkar et al. (2018); Kamath et al. (2020). We draw motivation from methods for cost-sensitive learning Mienye and Sun (2021) and selective classification Geifman and El-Yaniv (2017), which navigate penalties for failure. These settings reward confidence calibration, which can be critical for effective collaboration with human experts Verma et al. (2023). We are the first to investigate how serialized test-time compute helps models identify when they should not answer.
Figure 5: Jeopardy utility scales differently across models and thresholds. Performance of S1-32B and R1-32B in the Jeopardy odds setting under different confidence thresholds. Although S1 is competitive at lower budgets when the confidence threshold is 0, a higher threshold shows R1’s superior scaling performance in a selective setting.
5.1 Test-time Scaling
Many methods for scaling test-time compute have been explored. These include searching over possible generations (Wang et al., 2024), sampling many completions and selecting the best answer among them Wu et al. (2025), making gradient updates at inference time Akyürek et al. (2024); Li et al. (2025), using reinforcement learning to incentivize generating chains of thought before answering DeepSeek-AI et al. (2025), and simply fine-tuning on longer chains of thought (Muennighoff et al., 2025). Our work considers models fine-tuned on very long reasoning chains, and augments them with the ability to refuse to answer questions where they lack confidence. Concurrent work finds that such models can learn to predict when their answers are unlikely to be correct Zhang et al. (2025); Huang et al. (2025), but do not show how this affects performance as additional compute is expended. In contrast, we show how model confidence scales with test-time compute, and demonstrate its value for optimizing performance in settings that allow refusal to answer.
5.2 Selective Question Answering
Refusing to answer is an important option in many prior works on question answering. SQuAD 2.0 included this feature by asking questions which have no answer, although they treat abstaining from answering as a correct answer in these unanswerable cases Rajpurkar et al. (2018). Game-show based research efforts use an approach more closely aligned with ours, which penalizes systems for answering incorrectly to encourage abstentions when a system cannot develop sufficiently high confidence Ferrucci et al. (2010); Ferrucci (2012); Boyd-Graber and Börschinger (2020); Rodriguez et al. (2021). Related to quiz game settings is research into selective classification, which evaluates models performance across the coverage-accuracy curve, rather than at single point Geifman and El-Yaniv (2017). These approaches can be useful for avoiding costly errors in high-pressure domains Khan et al. (2018), under distribution shift Ren et al. (2023), or when designing systems that defer to expert humans when it lacks confidence that its input will be helpful Mozannar and Sontag (2020). Recent research in language modeling has investigated training language models to refuse to answer Cao (2024), and this capacity for refusal has become a point of competition among top industrial labs Wei et al. (2024). However, this line of work does not investigate this behavior in sequential test-time scaling models on reasoning intensive tasks, where a model might find a confident answer given higher compute budgets.
@@ 6 Conclusion
The selection function we implement is based entirely on the likelihood that a large language model assigns a series of tokens after thinking, which is not necessarily the optimal method for model confidence estimation. Furthermore, the Chain-of-Thought scaling method we apply may struggle to generalize to problem types that a model has note The method we use for ‘budget forcing’ (Muennighoff et al., 2025) may diminish performance by abruptly truncating chains of thought and driving the model outside of its training distribution. Concurrent work has introduced more elegant forms of compute budget control Aggarwal and Welleck (2025); Hou et al. (2025). Furthermore, we do not consider how compute costs might be incorporated in the model’s utility function, which could encourage increased energy consumption. Finally, we recognize that by evaluating only on English questions and answers, we may miss model capabilities or weaknesses in lower-resource languages or in multilingual settings.
Acknowledgments
This work was funded in part by the U.S. National Science Foundation under grant No. 2204926, and by the Defense Advanced Research Project Agency (DARPA) SciFy program (contract No. HR001125C0304) and CODORD program. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or DARPA. We thank Miriam Wanner, Zhengping Jiang, Orion Weller, Marc Marone, Alex Martin, and Beepul Bharti for helpful conversations throughout this project.
@@ Appendix A Appendix
Figure 6: (above) S1-32B’s answer accuracy is a function of compute budget and confidence threshold. This plot corresponds to the R1-32B plot in Figure 1. (below) Utility surface of S1-32B for Jeopardy. This plot corresponds to the R1-32B plot in Figure 4.
@@
Appendix B GPQA Experiments
GPQA consists of 448 graduate level multiple choice questions in the fields of physics, biology, and chemistry. These questions are considered “Google-proof”, in that skilled non-experts with access to the open internet struggle to answer them. We run experiments on the ‘Diamond’ subset of this dataset, which consists of the 198 questions which had the clearest answers to domain experts, while being the most difficult for non-experts to answer; this subset has also served as a benchmark for test-time scaling language models Muennighoff et al. (2025). Our experiments on GPQA follow the same basic procedure described in Section 3, except that we stop evaluate token budgets in range [500, 4000].
As with AIME, Figure 8 shows how models can increase performance by only answering when highly confident. Notably, there is very little gap between thresholds 0 and 0.5 on GPQA, owing to its multiple choice format. Moreover, whereas R1-32B outperformed S1-32B at all confidence thresholds, R1-32B only exceeds S1-32B at a confidence threshold of 0.95.
Figure 9 shows performance at Jeopardy odds of R1-32B (above) and S1-32B (below). Although both models perform similarly at low thresholds, R1-32B’s superior test-time scaling behavior becomes evident at higher confidence thresholds, allowing it to achieve higher peak utilities compared to S1-32B. This gap is reflected in the slices shown in Figure 10. We also note that the slices of the surface are nearly identical at lower confidence thresholds below 0.25. This is likely due to the multiple choice format of GPQA; when the model does not know the answer to a question, it assigns roughly equal probability to the four possible answers.
Figure 8: Confidence thresholds on test-time scaling. (left) When the confidence threshold is 0, the model answers 100% of questions. This is the only performance curve that is reported by test-time scaling research. (center) At a moderate threshold, more frequent absentions allow higher response accuracy. (right) At a high threshold, small amounts of test-time compute deliver very high accuracy, while test-time scaling provides more answers at the cost of answer accuracy. We treat the decision to never answer as yielding accuracy 0.
Figure 9: Utility surfaces of R1-32B (above) and S1-32B (below) for Jeopardy utility on GPQA. Utility is a function of compute budget and confidence threshold. These plots mirror the surfaces in Figure 4 and Figure 6.
Figure 10: Jeopardy utility on GPQA scales differently across models and thresholds. Performance of S1-32B and R1-32B on GPQA in the Jeopardy odds setting under different confidence thresholds. While S1 is competitive in the case when threshold is 0, a higher threshold shows R1’s superior scaling performance.